OpenStudy (anonymous):

Find the sum of the infinte geometric series: 24 + 12 + 6 + 3 + ...

OpenStudy (anonymous):

3/2

OpenStudy (anonymous):

how

OpenStudy (anonymous):

it keeps dividing by 2

OpenStudy (anonymous):

or keeps getting cut in half....well half of 3 is 1.5 or 3/2

OpenStudy (anonymous):

yes but im not looking for what is next im looking for the sum

OpenStudy (anonymous):

wait hold on

OpenStudy (anonymous):

you have to use a formula

OpenStudy (anonymous):

it says "find the sum" right?

OpenStudy (anonymous):

okay...i misunderstood the question apparently

OpenStudy (anonymous):

yes find the sum

OpenStudy (anonymous):

\[a+ar+ar^2+ar^3+...=\frac{a}{1-r}\] and here \[a=24,r=\frac{1}{2}\]

OpenStudy (anonymous):

sat...would it not be n/2?

OpenStudy (anonymous):

so you get \[\frac{24}{1-\frac{1}{2}}=\frac{24}{\frac{1}{2}}=24\times 2=48\]

OpenStudy (anonymous):

yours certainly looks better...i'm goin back to bed

OpenStudy (anonymous):

steps clear or no?

OpenStudy (anonymous):

yes thanks so much can i ask you another question

OpenStudy (anonymous):

a(1-r^n)/(1-r)

OpenStudy (anonymous):

derived

OpenStudy (anonymous):

sure post away, but make a new thread

OpenStudy (anonymous):

becomes

OpenStudy (anonymous):

oh u did it already crap

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

my connection slow

OpenStudy (anonymous):

its ok but thanks anyway :)